A revenue function is R(x)=22x and a cost function is C(x)= -9x + 341. The break-even point is:
a. (17, 374)
b. (242, 9)
c. (11, 242)
d. (31.5, 661.5)
: A revenue function is R(x)=22x and a cost function is C(x)= -9x + 341. The break-even point is:
a. (17, 374)
b. (242, 9)
c. (11, 242)
d. (31.5, 661.5)

Answer by nabla(410)

We want R=C.
So
22x=-9x+341
0=-31x+341
0=-31(x-11)
So x must=11.

R(11)=22*11=242.
So, (11,242).